### v2

# Observer Design over Hypercomplex Quaternions

Abstract
We develop observer design over hypercomplex quaternions in a characteristic-polynomial-free framework. Using the standard right-module convention, we derive a right observable companion form and its companion polynomial that encodes error dynamics via right-eigenvalue similarity classes. The design mirrors the real/complex case—coefficient updates in companion coordinates, followed by a similarity back—yet avoids determinants, characteristic/minimal polynomials, and Cayley–Hamilton identities that do not transfer to quaternions. We also give an Ackermann-type construction for the important case of closed-loop companion polynomials with real coefficients, ensuring similarity-equivariant evaluation. The results yield simple recipes for full-order observers directly over quaternions, clarify the role of right spectra and their similarity classes, and pinpoint when classical one-shot formulas remain valid. Numerical examples illustrate the method and advantages over vectorized or complex-adjoint surrogates.
I INTRODUCTION
Quaternions extend $\mathbb{R}$ and $\mathbb{C}$ to the division algebra $\mathbb{H}$, but—unlike $\mathbb{R}$ and $\mathbb{C}$—multiplication in $\mathbb{H}$ is noncommutative. This single feature alters standard observer machinery: right and left eigenvalues must be distinguished, and (on the right) eigenvalues are defined only up to quaternionic similarity classes. Familiar scalar proxies—determinants, characteristic polynomials, a full Cayley–Hamilton theorem—do not carry over verbatim; even polynomial evaluation depends on whether coefficients act on the left or on the right, so left and right zeros need not coincide. As a result, several classical constructions require new formulations when one works natively over $\mathbb{H}$; see, e.g., quaternionic stability [16] and linear algebra over $\mathbb{H}$ [18].
In this paper, we develop an observer-focused, determinant–free framework built on observability and its duality with controllability. We construct the observable companion (observability-form) realization directly from the observability matrix, link its right spectrum to companion-polynomial zeros, and derive a quaternionic Observer Ackermann formula that assigns right–eigenvalue similarity classes using a desired real-coefficient polynomial. Throughout, we remain in the hypercomplex (scalar) model of $\mathbb{H}$ where signals, states, and operators live directly in $\mathbb{H}.$ We work with the right spectra, and use left substitution for the companion polynomial—choices that make noncommutativity explicit and usable.
In aerospace and robotics applications, by contrast, quaternions are typically treated in their real $4$-vector (scalar–vector) form, and observers/filters are implemented with purely real-vector operations and models [13]. These methods are highly effective for attitude estimation and remain the default in those domains.
Here, however, we adopt the hypercomplex definition $q\in\mathbb{H}$ and work directly with quaternionic states and operators. This preserves the noncommutative algebra essential for right/left spectra and similarity classes, for companion/observability forms, and for a central (real-coefficient) polynomial calculus $a(A)$ without coordinate artifacts. The viewpoint is also motivated by areas where the hypercomplex structure is intrinsic—e.g., quantum dynamics and control [5, 12], NMR/MRI pulse design [6], polarization/bivariate signal processing and color imaging [10], quaternion-valued filters [20], quantum computing [14] and emerging quaternionic learning models [19, 15]. We believe the observer results established here over $\mathbb{H}$—including the observable companion construction and the quaternionic Observer Ackermann formula—are original.
II LINEAR ALGEBRA OVER QUATERNIONS
Quaternions
A quaternion is a hypercomplex scalar $q=a+b\mathbf{i}+c\mathbf{j}+d\mathbf{k}$ with $a,b,c,d\in\mathbb{R}$ and units $\mathbf{i},\mathbf{j},\mathbf{k}$ satisfying $\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=\mathbf{i}\mathbf{j}\mathbf{k}=-1$. The set $\mathbb{H}$ is a four-dimensional associative but noncommutative division algebra over $\mathbb{R}$. Write $\operatorname{Re}(q)=a$, $\operatorname{Im}(q)=b\mathbf{i}+c\mathbf{j}+d\mathbf{k}$, conjugate $\bar{q}=a-b\mathbf{i}-c\mathbf{j}-d\mathbf{k}$, and norm $|q|=\sqrt{q\bar{q}}$.
Other coordinate parameterizations are common. In the scalar–vector form, a quaternion is the real pair $q_{\mathrm{sv}}=(v_{\mathrm{s}},v_{\mathrm{v}})$, where the scalar part is $v_{\mathrm{s}}=a$ and the vector part is $v_{\mathrm{v}}=[\,b,\;c,\;d\,]^{\mathsf{T}}\in\mathbb{R}^{3}$. Equivalently, one stacks these into the vector–scalar $4$-vector $q_{\mathrm{sv}}=\begin{bmatrix}v_{\mathrm{v}}^{\mathsf{T}}&v_{\mathrm{s}}\end{bmatrix}^{\mathsf{T}}\in\mathbb{R}^{4}.$ These representations are standard in aerospace and robotics.
In this paper, we work natively over $\mathbb{H}$ to preserve the noncommutative structure needed for right/left spectra and observable–companion constructions. For details, see [18].
Similarity classes
Quaternions $q,r$ are similar, denoted $q\sim r$, if $q=\alpha^{-1}r\alpha$ for some nonzero $\alpha\in\mathbb{H}$. Similarity preserves $\operatorname{Re}(\cdot)$, $|\cdot|$, and $|\operatorname{Im}(\cdot)|$ and partitions $\mathbb{H}$ into equivalence classes $\cdot$]. Real quaternions are isolated: $a]=\{a\}$. If $\operatorname{Im}(q)\neq 0$, then $q]=\{\operatorname{Re}(q)+u|\operatorname{Im}(q)|:u\in\mathrm{span}\{\mathbf{i},\mathbf{j},\mathbf{k}\},|u|=1\}$, a 2-sphere meeting the standard complex line $\{a+b\mathbf{i}:a,b\in\mathbb{R}\}$ in the two points $\operatorname{Re}(q)\pm\mathbf{i}|\operatorname{Im}(q)|$. For instance, $\mathbf{k}]$ contains all imaginary units, including $\pm\mathbf{i}$.
Matrices over $\mathbb{H}$
For $m,n\in\mathbb{N}$, let $\mathbb{H}^{m\times n}$ denote the set of $m\times n$ matrices with entries in $\mathbb{H}.$ We write $A^{*}=\bar{A}^{\mathsf{T}}$ (entry-wise conjugation followed by transpose). Noncommutativity affects basic identities: in general $\overline{AB}\neq\bar{A}\bar{B}$ and $(AB)^{\mathsf{T}}\neq B^{\mathsf{T}}A^{\mathsf{T}}$, while the mixed rule $(AB)^{*}=B^{*}A^{*}$ does hold. See [18] for further details.
Caveat for design. Classical $\mathbb{R}/\mathbb{C}$ tools—determinant, characteristic polynomial, and the Cayley–Hamilton theorem—do not carry over verbatim to $\mathbb{H}$. The determinant admits nonstandard variants with different meanings; a “characteristic polynomial” in the usual sense may be absent or enjoy only limited properties; and Cayley–Hamilton can fail or requires alternative formulations. Hence these tools cannot be used naively for observer design over $\mathbb{H}$ and are replaced here by determinant-free, characteristic-polynomial-free constructions.
Rank and invertibility
For $A\in\mathbb{H}^{m\times n}$, the quaternionic rank $\operatorname{rank}_{\mathbb{H}}A$ is the maximal number of right–independent columns (i.e., if $\sum_{j}a_{j}c_{j}=0$ with $c_{j}\in\mathbb{H}$ on the right, then all $c_{j}=0$). A square $A\in\mathbb{H}^{n\times n}$ is invertible iff $\operatorname{rank}_{\mathbb{H}}A=n$ [18].
Right eigenvalues and the spectrum
For $A\in\mathbb{H}^{n\times n}$, a right eigenpair $(v,\lambda)$ satisfies
$Av$ $=$ $v\lambda,\qquad v\in\mathbb{H}^{n}\setminus\{0\},\ \lambda\in\mathbb{H}.$ (1)
Right scaling of $v$ by $\alpha\neq 0$ replaces $\lambda$ by $\alpha^{-1}\lambda\alpha$, so right eigenvalues are defined only up to similarity. The right spectrum is thus the set of classes $\sigma_{\mathscr{r}}(A):=\{[\lambda]:\exists v\neq 0\text{ with }Av=v\lambda\}\subset\mathbb{H}/\!\sim,$ which is invariant under similarity $A\mapsto S^{-1}AS$ [18]. When needed, take the unique complex representative with nonnegative imaginary part—the standard eigenvalue.
Polynomials and evaluation
Let $\mathbb{H}[\lambda]$ be the ring of quaternionic polynomials $p(\lambda)=\sum_{k=0}^{n}p_{k}\lambda^{k}$ with $p_{k}\in\mathbb{H}$ and indeterminate $\lambda$ commuting with all coefficients. Evaluation at $q\in\mathbb{H}$ depends on side: the right value is $p(q)_{\mathscr{r}}=\sum_{k=0}^{n}p_{k}q^{k}$, while the left value is $p(q)_{\ell}=\sum_{k=0}^{n}q^{k}p_{k}$.
Left/right zeros
A right zero of $p$ is $q$ with $p(q)_{\mathscr{r}}=0$; a left zero satisfies $p(q)_{\ell}=0$. Left and right zeros correspond up to similarity: if $p(q)_{\ell}=0$, then for some $\alpha\neq 0$, $p(\alpha^{-1}q\alpha)_{\mathscr{r}}=0$, and conversely [8]. If all $p_{k}\in\mathbb{R}$ (central), then $p(q)_{\mathscr{r}}=p(q)_{\ell}$ and the two zero sets coincide.
Isolated vs. spherical zeros
A degree-$n$ quaternionic polynomial has either $n$ (counting multiplicity) right zeros or infinitely many. A nonreal zero is either isolated (a single point in its class) or spherical (the full class $[q]$, a 2-sphere). If the coefficients are real, every zero is either real (hence isolated) or spherical [17].
Notation
We regard $\mathbb{H}^{n}$ as a right $\mathbb{H}$-module. Let $e_{i}$ be the $i$th canonical basis vector of $\mathbb{R}^{n}\subset\mathbb{H}^{n}$ (a $1$ in position $i$, zeros elsewhere); in particular, $e_{n}=[0,\ldots,0,1]^{\mathsf{T}}$. Because these entries are real (central), left/right scalar multiplication is unambiguous whenever $e_{i}$ appears.
III QUATERNIONIC STATE–SPACE SISO SYSTEMS
We study continuous-time linear time-invariant (LTI) single-input single-output (SISO) systems over the division algebra $\mathbb{H}$.
Model
In state-space form, the dynamics are
$\dot{x}(t)$ $=Ax(t)+Bu(t),$ (2)
$y(t)$ $=Cx(t)+Du(t),$
where $A\in\mathbb{H}^{n\times n}$, $B\in\mathbb{H}^{n}$ (column), $C\in\mathbb{H}^{1\times n}$ (row), $D\in\mathbb{H}$, the state vector is $x(t)\in\mathbb{H}^{n}$, and the input/output scalars are $u(t),y(t)\in\mathbb{H}$.
Conventions
We adopt the right-module convention: matrices act from the left and quaternionic scalars multiply signals on the right. Thus the products $Bu(t)$ and $Du(t)$ are right-multiplications by $u(t)$; this guarantees right-linearity in both $x$ and $u$.
Internal stability
With $u\!\equiv\!0$, stability means $x(t)\!\to\!0$ (CT) or $x(k)\!\to\!0$ (DT) for all initial states. Equivalently, using the standard right eigenvalues $\lambda$ of $A$, the system is stable iff $\operatorname{Re}\lambda<0$ for all $\lambda$; (see [16]).
Controllability
As in the real/complex case, controllability means the ability to steer the state from any $x_{0}\in\mathbb{H}^{n}$ to any $x_{f}\in\mathbb{H}^{n}$ within a finite horizon by a suitable input. In the single–input setting, the controllability matrix is
$\mathcal{C}$ $=$ $\bigl[BAB\cdots A^{n-1}B\bigr]\in\mathbb{H}^{n\times n},$
and the pair $(A,B)$ is controllable iff $\mathcal{C}$ is invertible over $\mathbb{H}$ (equivalently, iff $\operatorname{rank}_{\mathbb{H}}\mathcal{C}=n$) [9, 18, 21].
Observability
Dually, observability means that the initial state can be uniquely reconstructed from finitely many output measurements. In the single–output setting, the observability matrix is
$\mathcal{O}$ $=$ $\begin{bmatrix}C\\ CA\\ \vdots\\ CA^{n-1}\end{bmatrix}\in\mathbb{H}^{n\times n},$ (3)
and the pair $(A,C)$ is observable iff $\mathcal{O}$ is invertible over $\mathbb{H}$ (equivalently, iff $\operatorname{rank}_{\mathbb{H}}\mathcal{O}=n$) [9, 18, 21].
Similarity
All rank conditions are over $\mathbb{H}$ and are invariant under similarity: $A\mapsto S^{-1}AS$, $B\mapsto S^{-1}B$, $C\mapsto CS$ for any invertible $S\!\in\!\mathbb{H}^{n\times n}$.
IV OBSERVABLE COMPANION FORM AND COMPANION POLYNOMIAL
In the classical (real/complex) LTI setting, two staple tools are the observable companion form and the characteristic or minimal polynomial. For quaternionic models, the observable companion form remains valid. By contrast, characteristic and minimal polynomials do not translate verbatim, since quaternions lack a straightforward determinant. Consequently, we replace them with a suitably defined quaternionic companion polynomial.
Definition 1 (Observable companion form and companion polynomial).
Let $A_{o}\in\mathbb{H}^{n\times n}$ and $C_{o}\in\mathbb{H}^{1\times n}$ be
$A_{o}$ $=$ $\begin{bmatrix}0&\ldots&0&-a_{0}\\ 1&\ldots&0&-a_{1}\\ \vdots&\ddots&\vdots&\vdots\\ 0&\ldots&1&-a_{n-1}\end{bmatrix}$, (4)
$C_{o}$ $=$ $\begin{bmatrix}0&0&\ldots&1\end{bmatrix}$, (5)
with $a_{0},a_{1},\ldots,a_{n-1}\in\mathbb{H}$. We say that $(A_{o},C_{o})$ is in right observable companion (canonical) form. The companion polynomial associated with $(A_{o},C_{o})$ (or with $A_{o}$) is the monic quaternionic polynomial
$a(\lambda)$ $=$ $a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{H}[\lambda].$
If $A\in\mathbb{H}^{n\times n}$ is cyclic with respect to the row CC—i.e., there exists $C\in\mathbb{H}^{1\times n}$ with the observability matrix $\mathcal{O}$ invertible—then $A$ is similar to the right observable companion form. The next theorem gives a constructive procedure.
Theorem 1 (Determinant-free companion form).
Let $A\in\mathbb{H}^{n\times n}$ and $C\in\mathbb{H}^{1\times n}$. Assume the SISO observability matrix $\mathcal{O}$ from (3) is invertible over $\mathbb{H}$.
  1. 1.
    There exists an invertible similarity $T\in\mathbb{H}^{n\times n}$ such that
    $A_{o}\;=\;T^{-1}AT,\qquad C_{o}\;=\;CT,$
    and the pair $(A_{o},C_{o})$ is in the right observable companion form
[…]
    where the scalars $a_{0},\ldots,a_{n-1}\in\mathbb{H}$ are defined as the (negatives of the) entries in the rightmost column of $A_{o}$.
  2. 2.
    The corresponding similarity matrix is uniquely determined as
    $s$ $:=$ $\mathcal{O}^{-1}e_{n}\ \in\ \mathbb{H}^{n},$
    $T$ $:=$ $\begin{bmatrix}s&As&\cdots&A^{n-1}s\end{bmatrix}$, (7)
  3. 3.
    For the given pair $(A,C)$, the matrices $A_{o},C_{o}$ constructed via (7) and the coefficients $a_{k}$ (as defined above) are unique.
Proof:
We adapt the classical real/complex argument (cf. [2]) to the quaternionic setting, avoiding determinants and Cayley–Hamilton.
Choice of basis and similarity. Let $s:=\mathcal{O}^{-1}e_{n}$ and $T=[\,s\ As\ \cdots\ A^{n-1}s\,]$. From $\mathcal{O}s=e_{n}$ we obtain $Cs=\cdots=CA^{n-2}s=0$ and $CA^{n-1}s=1$, hence
$CT=\bigl[\,Cs\ \ CAs\ \ \cdots\ \ CA^{n-1}s\,\bigr]=e_{n}^{\mathsf{T}}.$
Moreover, for each $k=0,\ldots,n-1$, the row $CA^{k}T$ has the triangular pattern
$CA^{k}T=\bigl[\,\underbrace{0\ \cdots\ 0}_{n-k-1}\ \ 1\ \ \star\ \cdots\ \star\,\bigr],$ (8)
where $\star$’ denotes unspecified entries. Hence, $(CA^{k}T)e_{j}=0$ for j<nkj<n-k and $(CA^{k}T)e_{n-k}=1$.
Define the transformed pair. Set $M:=T^{-1}AT$ and $C_{o}:=CT=e_{n}^{\mathsf{T}}$.
Output row. From $CT=e_{n}^{\mathsf{T}}$ we have $C_{o}=e_{n}^{\mathsf{T}}=[\,0\ \cdots\ 0\ 1\,]$.
Subdiagonal shift property. Write $T=[w_{1},\ldots,w_{n}]$ with $w_{1}=s$ and $w_{j+1}=Aw_{j}$ by construction. Then
$Me_{j}=T^{-1}ATe_{j}=T^{-1}Aw_{j}=T^{-1}w_{j+1}=e_{j+1},$
for $j=1,\ldots,n-1,$ so $M$ has ones on the subdiagonal and zeros below it.
Rightmost column and the definition of $a_{k}$. By definition of the observable companion form, the last column of $M$ is some vector $a_{0},\ldots,a_{n-1}]^{\mathsf{T}}\in\mathbb{H}^{n}$. We define the coefficients $a_{k}\in\mathbb{H}$ to be these entries. With the output row and subdiagonal shift established above, this uniquely identifies $M$ as the observable companion $A_{o}$ and yields (6).
Uniqueness. The vector $s$ is uniquely determined by $\mathcal{O}s=e_{n}$, hence $s=\mathcal{O}^{-1}e_{n}$. Then the recursion $w_{j+1}=Aw_{j}$ uniquely fixes the columns of $T=[\,s,As,\ldots,A^{n-1}s\,]$. With $C_{o}=e_{n}^{\mathsf{T}}$, subdiagonal ones and the last column $a_{0},\ldots,a_{n-1}]^{\mathsf{T}}$, the matrix $A_{o}=M$ is unique as well. ∎
Remark 1.
In the real/complex (central) case, the last-column coefficients $a_{k}$ also satisfy $CA^{n}=-\sum_{k=0}^{n-1}(CA^{k})\,a_{k}$. Over $\mathbb{H}$, a right–linear expansion still holds but with (generally different) scalars $b_{k}\in\mathbb{H}$: $CA^{n}=-\sum_{k=0}^{n-1}(CA^{k})\,b_{k}$, and there is no simple closed-form relation between $a_{k}$ and $b_{k}$ due to noncommutativity; if the $a_{k}$ are real, then $b_{k}=a_{k}$.
Remark 2.
Compared with the classical $\mathbb{R}/\mathbb{C}$ case, the construction proceeds in reverse. Instead of first computing a characteristic or minimal polynomial (via determinants) and then forming a companion realization, we directly build the observable companion form from the invertible observability matrix $\mathcal{O}$, and only afterwards read off the coefficients of the companion polynomial from the rightmost column of $A_{o}$. This approach is determinant–free and avoids the use of characteristic polynomials, Cayley–Hamilton, or adjugates, whose straightforward generalizations to $\mathbb{H}$ are problematic.
Example 1.
Let
$A$ $=$ $\tfrac{1}{4}\begin{bmatrix}-2+\mathbf{j}&\mathbf{i}\\ \mathbf{i}&-2-\mathbf{j}\end{bmatrix},\;\;C=\begin{bmatrix}\mathbf{j}&\mathbf{k}\end{bmatrix}$. (9)
Then the observability matrix inverse is
$\mathcal{O}^{-1}=\tfrac{1}{2}\begin{bmatrix}-1-3\mathbf{j}&-4\mathbf{j}\\ \mathbf{i}+\mathbf{k}&4\mathbf{k}\end{bmatrix}$. (10)
and the similarity transformation matrix (7) reads
$T=\tfrac{1}{2}\begin{bmatrix}-4\mathbf{j}&1+\mathbf{j}\\ 4\mathbf{k}&-\mathbf{i}-3\mathbf{k}\end{bmatrix}$. (11)
The right observable companion form of $A$ is
$A_{o}$ $=$ $\begin{bmatrix}0&-0.25+0.25\mathbf{j}\\ 1&-1+0.5\mathbf{j}\end{bmatrix},\;C_{o}=\begin{bmatrix}0&1\end{bmatrix}$, (12)
and the associated companion polynomial is
$a(\lambda)$ $=$ $0.25-0.25\mathbf{j}+(1-0.5\mathbf{j})\lambda+\lambda^{2}.$ (13)
We obtain an equally clean annihilating identity for the right observable companion, again without invoking a full Cayley–Hamilton theorem over $\mathbb{H}$ (cf. [3, 18]). We will use left substitution for the companion polynomial at $A_{o}.$
Theorem 2 (Companion polynomial annihilates its observable companion matrix over $\mathbb{H}$).
Let $A_{o}\in\mathbb{H}^{n\times n}$ be in observable companion form
$A_{o}$ $=$ $\begin{bmatrix}0&0&\cdots&-a_{0}\\ 1&0&\cdots&-a_{1}\\ \vdots&\ddots&\ddots&\vdots\\ 0&\cdots&1&-a_{n-1}\end{bmatrix}$, (14)
with $a_{0},\ldots,a_{n-1}\in\mathbb{H}$, and define the companion polynomial $a(\lambda)=a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}\ \in\mathbb{H}[\lambda].$ Then $a(\lambda)$ annihilates $A_{o}$ under left substitution:
$Ia_{0}+A_{o}a_{1}+\cdots+A_{o}^{n-1}a_{n-1}+A_{o}^{n}=0.$ (15)
Proof:
Setup and evaluation convention. Let $A_{o}$ be in the observable companion form (14) with last column $\bigl[a_{0},\ldots,a_{n-1}\bigr]^{\mathsf{T}}$, where $a_{k}\in\mathbb{H}$. Following Section 2, we evaluate polynomials at $A_{o}$ by left substitution:
$a(A_{o})_{\ell}=A_{o}^{n}+\sum_{k=0}^{n-1}A_{o}^{k}\,a_{k}.$ (16)
Our goal is to show $a(A_{o})_{\ell}=0$.
Krylov columns and the shift property. Define $u_{k}:=A_{o}^{k}e_{1}$ for $k\geq 0$. Since $A_{o}$ has ones on the subdiagonal,
$A_{o}e_{j}=e_{j+1}\quad(j=1,\ldots,n-1),$
hence $u_{k}=e_{k+1}\,\;(k=0,\ldots,n-1)$, and thus
$\mathcal{K}:=\begin{bmatrix}u_{0}&\ldots&u_{n-1}\end{bmatrix}=I_{n}.$ (17)
The step-$n$ recurrence from the last column. Reading off the last column of $A_{o}$ gives
$A_{o}^{n}e_{1}=-\begin{bmatrix}a_{0}\\ a_{1}\\ \vdots\\ a_{n-1}\end{bmatrix}=-\sum_{k=0}^{n-1}u_{k}\,a_{k}.$
Left-multiplying by $A_{o}^{j}$ and using $A_{o}^{j}u_{k}=u_{k+j}$ yields, for all $j\geq 0$,
$u_{n+j}=-\sum_{k=0}^{n-1}u_{k+j}\,a_{k}.$ (18)
Annihilation under left substitution. For any $j\geq 0$,
$a(A_{o})_{\ell}\,u_{j}$ $=$ $\Bigl(A_{o}^{n}+\sum_{k=0}^{n-1}A_{o}^{k}a_{k}\Bigr)u_{j}$ (19)
$=$ $u_{n+j}+\sum_{k=0}^{n-1}u_{k+j}\,a_{k}=0$
by (18). In particular, this holds for $j=0,1,\ldots,n-1$, so
$a(A_{o})_{\ell}\;\mathcal{K}=0.$
Since $\mathcal{K}=I_{n}$, we conclude $a(A_{o})_{\ell}=0$, which is the claimed annihilation under left substitution. ∎
Remark 3.
This is not a quaternionic Cayley–Hamilton theorem (which generally fails over $\mathbb{H}$). Rather, it is a one-sided annihilation under left substitution that hinges on the observable companion’s shift/subdiagonal structure and the rightmost-column read-off of the coefficients. It neither asserts a two-sided polynomial identity nor identifies $a(\lambda)$ with any characteristic/minimal polynomial; it also does not automatically extend to arbitrary similar realizations outside the observable companion form.
The right spectrum of the observable companion matrix is likewise composed of the similarity classes of the right zeros of its companion polynomial.
Theorem 3 (Right zeros form the right spectrum of the observable companion matrix).
Let $A_{o}\in\mathbb{H}^{n\times n}$ be in observable companion form (14), and let $a(\lambda)=a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{H}[\lambda]$ be its companion polynomial (with $a_{k}\in\mathbb{H}$ defined by the last column of $A_{o}$). Then:
  1. 1.
    If $\lambda\in\mathbb{H}$ is a right zero of $a(\lambda)$, then $\lambda$ is a right eigenvalue of $A_{o}$. The corresponding eigenvector $v=\begin{bmatrix}v_{1},\ldots,v_{n}\end{bmatrix}$ is obtained by the backward recurrence
    $v_{n}=\!1;\;v_{j-1}\!=\!a_{j-1}\!+\!v_{j}\lambda,\;j=n,n-1,\ldots,2,$ (20)
    and satisfies $A_{o}v=v\lambda$.
  2. 2.
    Hence, the right spectrum of $A_{o}$ is exactly the union of the similarity classes of the right zeros of $a(\lambda)$:
    $\sigma_{\mathscr{r}}(A_{o})=\mathbin{{\cup}}_{i}[\lambda_{i}],$
    counting multiplicities.
  3. 3.
    Similarly, every left zero $\mu$ of $a(\lambda)$ is (up to similarity) also a right eigenvalue of $A_{o}$.
Proof:
(1) Right zeros $\Rightarrow$ right eigenpairs. Assume $a(\lambda)_{\mathscr{r}}=0$ and construct $v\in\mathbb{H}^{n}$ by the backward recurrence (20). We claim $A_{o}v=v\lambda$.
For rows $j=2,\ldots,n$ of $A_{o}v$, the observable companion structure gives
$(A_{o}v)_{j}$ $=$ $v_{j-1}-a_{j-1}\,v_{n}$
$=$ $(v_{j}\lambda+a_{j-1})-a_{j-1}=v_{j}\lambda=(v\lambda)_{j}.$
For the first row,
$(A_{o}v)_{1}$ $=$ $-a_{0}\,v_{n}\;=\;-a_{0},$
$(v\lambda)_{1}$ $=$ $v_{1}\lambda=\bigl(a_{1}+\cdots+a_{n-1}\lambda^{n-2}+\lambda^{n-1}\bigr)\lambda$
$=$ $a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}.$
Hence $(A_{o}v)_{1}=(v\lambda)_{1}$ is equivalent to $a_{0}+\sum_{k=1}^{n-1}a_{k}\lambda^{k}+\lambda^{n}=0$, i.e. $a(\lambda)_{\mathscr{r}}=0$, which holds by assumption. Therefore $A_{o}v=v\lambda$ and $\lambda$ is a right eigenvalue.
(2) Right eigenpairs $\Rightarrow$ zeros up to similarity. Conversely, suppose $A_{o}x=x\mu$ with $x\neq 0$. From rows $j=2,\ldots,n$ of the observable companion,
$x_{j-1}-a_{j-1}x_{n}=x_{j}\mu,$
which rearranges to $x_{j-1}=a_{j-1}x_{n}+x_{j}\mu$. Iterating backwards yields, for every $k=1,\ldots,n-1,$
$x_{k}=a_{k}x_{n}+a_{k+1}x_{n}\mu+\cdots+a_{n-k}x_{n}\mu^{n-k-1}+x_{n}\mu^{n-k}.$
In particular, for $k=1$
$x_{1}=a_{1}x_{n}+a_{2}x_{n}\mu+\cdots+a_{n-1}x_{n}\mu^{n-2}+x_{n}\mu^{n-1}.$
Using the first row of $A_{o}x=x\mu$ and substituting for $x_{1}$,
$-a_{0}x_{n}=x_{1}\mu=a_{1}x_{n}\mu+\cdots+a_{n-1}x_{n}\mu^{n-1}+x_{n}\mu^{n}$
Note that $x_{n}\neq 0$; otherwise, from the $n$-th row $x_{n-1}-a_{n-1}x_{n}=x_{n}\mu$ we get $x_{n-1}=0$, and iterating upward forces $x=0$, a contradiction. Hence set the similar quaternion
$\lambda:=x_{n}\mu x_{n}^{-1},$
so that $x_{n}\mu^{k}=\lambda^{k}x_{n}$ for all $k$. Then
$-a_{0}x_{n}=a_{1}\lambda x_{n}+\cdots+a_{n-1}\lambda^{n-1}x_{n}+\lambda^{n}x_{n},$
and right–cancelling $x_{n}$ gives
$a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}=0,$
i.e. $a(\lambda)_{\mathscr{r}}=0$. Thus every right eigenvalue $\mu$ is similar to a right zero $\lambda$, proving the spectral statement.
(3) Left zeros. By Gordon–Motzkin, left and right zeros of a quaternionic polynomial lie in the same similarity classes; hence every left zero is similar to a right zero and therefore (up to similarity) to a right eigenvalue of $A_{o}$.
Combining (1)–(3) gives the three claims of the theorem. ∎
Example 2.
For $A_{o}$ from (12) with $a(\lambda)$ from (13), we have
$a(A_{o})_{\ell}$ $=$ $I_{2}(0.25\!-\!0.25\mathbf{j})+\begin{bmatrix}0&\!-\!0.25\!+\!0.25\mathbf{j}\\ 1&\!-\!1\!+\!0.5\mathbf{j}\end{bmatrix}(1\!-\!0.5\mathbf{j})$
$\negmedspace{}+\begin{bmatrix}\!-\!0.25\!+\!0.25\mathbf{j}&0.125\!-\!0.375\mathbf{j}\\ \!-\!1\!+\!0.5\mathbf{j}&0.5\!-\!0.75\mathbf{j}\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$.
The companion polynomial can likewise be attached to any cyclic quaternionic matrix $A$ via its observable companion realization $(A_{o},C_{o})$.
Definition 2 (Companion polynomial of an observable pair).
Let $(A,C)$ with $A\in\mathbb{H}^{n\times n}$ and $C\in\mathbb{H}^{1\times n}$ be such that the observability matrix $\mathcal{O}$ from (3) is invertible over $\mathbb{H}$. Let $(A_{o},C_{o})$ be the unique right observable companion realization of $(A,C)$ (Theorem 1), and write the last column of $A_{o}$ as $\bigl[a_{0},\ldots,a_{n-1}\bigr]^{\mathsf{T}}$ with $a_{k}\in\mathbb{H}$. The companion polynomial of $(A,C)$ is the monic quaternionic polynomial
$a(\lambda)$ $=$ $a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}.$ (21)
Remark 4.
Throughout, we use the right observable companion form. Other orientations—left/upper/lower variants obtained by permutations—are equivalent and lead, after the obvious similarity/permutation, to the same monic companion polynomial.
Theorem 4 (Right zeros form the right spectrum of $A$ (observable case)).
Let $(A,C)$ be an observable quaternionic SISO pair and let $a(\lambda)$ be its companion polynomial from (21). Then
  1. 1.
    Any right root $\lambda\in\mathbb{H}$ of $a(\lambda)$ yields a right eigenvalue (up to similarity) of $A$.
  2. 2.
    Hence, the right spectrum of $A$ is exactly the union of the similarity classes of the right zeros of $a(\lambda)$: $\sigma_{\mathscr{r}}(A)=\mathbin{{\cup}}_{i}[\lambda_{i}]$, counting multiplicities.
  3. 3.
    Likewise, each left root $\mu$ of $a(\lambda)$ is similar to some right eigenvalue of $A$.
Proof:
By Theorem 1, AA is similar over $\mathbb{H}$ to its observable companion $A_{o}$, so $\sigma_{\mathscr{r}}(A)=\sigma_{\mathscr{r}}(A_{o})$. The claims then follow from Theorem 3. ∎
Remark 5.
As in the controllable case, the companion polynomial shares several roles of the characteristic/minimal polynomial but not all: in general, it does not annihilate $A$; It only annihilates the observable companion $A_{o}$. Hence $a(A_{o})_{\ell}=0$ while typically $a(A)_{\ell}\neq 0$.
Remark 6.
When $A$ is cyclic, it is natural to speak of the companion polynomial of $A$ constructed from any observable SISO pair $(A,C)$. Different cyclic choices of $C$ may yield different polynomials. Still, their right zeros agree up to quaternionic similarity classes: for every such $C$, the right zeros of the associated polynomial form the same multiset of classes—the right spectrum of $A$. For a fixed pair $(A,C)$, however, the observable companion $(A_{o},C_{o})$ and the attached polynomial are unique.
V OBSERVERS
An observer is a dynamic estimator that reconstructs the state $\hat{x}$ from known inputs $u$ and measured outputs $y$ by injecting the innovation $y-C\hat{x}$ through a gain $L$. It enables implementable state feedback $u=K\hat{x}$, monitoring, and prediction. The relations below are standard from the real/complex literature and will be referenced later. They also hold for quaternionic systems over $\mathbb{H}$ with minor modifications that are detailed in a later section.
A well-known Luenberger full-order observer [11] is described by
[…]
$e$ $:=$ $x-\hat{x},\quad\dot{e}=(A-LC)e.$
and the design goal is to place eigenvalues of the observer error dynamics (observer poles) $\sigma_{\mathscr{r}}(A-LC)$ in the open left half-plane.
VI DUALITY OVER $\mathbb{H}$
Over $\mathbb{R}$, transpose drives controller–observer duality: controllability of $(A,B)$ is equivalent to observability of $(A^{\mathsf{T}},B^{\mathsf{T}})$, and one often maps gains via $L=K^{\mathsf{T}}$ (cf. [2]).
Over $\mathbb{H}$, this fails because the correct adjoint is the conjugate transpose $M^{*}=\overline{M}^{\mathsf{T}}$, which alone preserves the quaternionic Hermitian form and converts left to right eigenpairs with eigenvalue conjugation [21, 18, 7, 1]. Thus, duality must be formulated with and with explicit care for right/left eigenpairs (see also quaternion-valued system treatments [9, 4]).
The original system and its dual system are related through mappings
$A\leftrightarrow A^{*},B\leftrightarrow C^{*},C\leftrightarrow B^{*},D\leftrightarrow D^{*}.$
Their controllability and observability are related via
$\mathcal{O}_{n}(A^{*},B^{*})$ $=$ $\mathcal{C}_{n}(A,B)^{*},$
see [21, 18], and therefore
$(A,B)\ \text{controllable}\ \Longleftrightarrow\ (A^{*},B^{*})\ \text{observable},$
$(A,C)\ \text{observable}\ \Longleftrightarrow\ (A^{*},C^{*})\ \text{controllable}.$
If $K\in\mathbb{H}^{1\times n}$ is a controller for the dual pair $(A^{*},C^{*})$, then the observer (output-injection) gain for the original system is
$L$ $=$ $K^{*}\in\mathbb{H}^{n\times 1}.$
The error and dual closed-loop matrices satisfy $\sigma_{\mathscr{r}}\big(A-LC\big)=\sigma_{\mathscr{r}}\big(A^{*}-C^{*}K\big).$ Since $q]=[\overline{q}]$ for quaternionic similarity classes, the two spectra coincide as class sets. See [7, 1].
Proof sketches and references
The transpose fails over $\mathbb{H}$; only the adjoint satisfies $\langle Ax,y\rangle=\langle x,A^{*}y\rangle$ and aligns with quaternionic spectral theory [21, 18, 7, 1]. The dual mapping $(A,B,C,D)\!\leftrightarrow\!(A^{*},C^{*},B^{*},D^{*})$ is standard. Using $(AB)^{*}=B^{*}A^{*}$ and rank invariance yields $\mathcal{O}_{n}(A^{*},B^{*})=\mathcal{C}_{n}(A,B)^{*}$, hence $(A,B)$ controllable $\Leftrightarrow$ $(A^{*},B^{*})$ observable and $(A,C)$ observable $\Leftrightarrow$ $(A^{*},C^{*})$ controllable [21, 18]. Design $K$ on $(A^{*},C^{*})$ and set $L=K^{*}$; then $\sigma_{\mathscr{r}}(A-LC)=\overline{\sigma_{\mathscr{r}}(A^{*}-C^{*}K)}=\sigma_{\mathscr{r}}(A^{*}-C^{*}K)$ by conjugation invariance of right-eigenvalue classes [7, 1].
VII QUATERNIONIC OBSERVER POLE PLACEMENT VIA COMPANION FORM
In $\mathbb{R}/\mathbb{C}$, observer design is dual to state feedback; over $\mathbb{H}$ the same SISO workflow applies: transform to observable companion form, read the companion polynomial from the rightmost column, interpret poles as right-eigenvalue similarity classes, update coefficients in observer coordinates, and map the gain back. The construction is determinant-free, avoiding characteristic/minimal polynomials, Cayley–Hamilton, and adjugates.
Theorem 5 (Pole placement via observable companion form).
Assume a SISO observable pair $(A,C)$ with $A\in\mathbb{H}^{n\times n}$, $C\in\mathbb{H}^{1\times n}$. Let $(A_{o},C_{o})$ be its right observable companion form via an invertible $T\in\mathbb{H}^{n\times n}$, i.e.
$A_{o}$ $=$ $T^{-1}AT,\quad C_{o}=CT=\bigl[0\ 0\ \cdots\ 1\bigr],$ (23)
and write the rightmost column of $A_{o}$ as $\bigl[-a_{0},\ldots,-a_{n-1}\bigr]^{\mathsf{T}}$. Let
$a(\lambda)$ $=$ $a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}$
be the associated companion polynomial. Given any monic
$a_{d}(\lambda)$ $=$ $d_{0}+d_{1}\lambda+\cdots+d_{n-1}\lambda^{n-1}+\lambda^{n},$
there exists a unique observer gain $L\in\mathbb{H}^{n\times 1}$ such that the right spectrum of
$A_{\mathrm{obs}}$ $:=$ $A-LC$ (24)
equals the multiset of right zeros (similarity classes, with multiplicities) of $a_{d}$. Moreover, with $L_{o}:=T^{-1}L$ we have
$L$ $=$ $TL_{o},$ (25)
$L_{o}$ $=$ $\bigl[d_{0}-a_{0}\ \ d_{1}-a_{1}\ \ \cdots\ \ d_{n-1}-a_{n-1}\bigr]^{\mathsf{T}}.$
Proof:
Work in observable coordinates. With $L_{o}:=T^{-1}L$ and (23),
$A_{\mathrm{obs},o}$ $:=$ $T^{-1}(A\!-\!LC)T\!=\!A_{o}\!-\!L_{o}C_{o}.$
Here $L_{o}C_{o}$ is the rank-one matrix whose last column equals $L_{o}$ and all other columns are zero. Writing $L_{o}=\bigl[\ell_{0},\ldots,\ell_{n-1}\bigr]^{\mathsf{T}}$, the rank–one update subtracts $L_{o}$ from the last (rightmost) column of $A_{o}$. The rightmost column becomes $\bigl[-(a_{0}+\ell_{0}),\dots,-(a_{n-1}+\ell_{n-1})\bigr]^{\mathsf{T}}$ which yields the coefficient update.
A direct computation of the companion polynomial of $A_{\mathrm{obs},o}$ shows that the closed–loop coefficients are updated as
$a_{\mathrm{obs}}(\lambda)$ $=$ $(a_{0}+\ell_{0})+(a_{1}+\ell_{1})\lambda+\cdots$
$\negmedspace{}+(a_{n-1}+\ell_{n-1})\lambda^{n-1}+\lambda^{n}.$
Choosing $\ell_{i}=d_{i}-a_{i}$ makes $a_{\mathrm{obs}}(\lambda)=a_{d}(\lambda)$. By the observable–companion root/spectrum property, the right zeros of $a_{\mathrm{obs}}$ are the right eigenvalues of $A_{\mathrm{obs},o}$, hence of $A_{\mathrm{obs}}$. Uniqueness follows since the coefficient update $\ell_{i}=d_{i}-a_{i}$ is bijective and $L\leftrightarrow L_{o}$ via $T$ is bijective. Finally (25) holds by definition. ∎
Example 3 (Example 1 continued).
With the companion form (12), the similarity matrix (11), and the companion polynomial (13) in hand, we now consider several choices of desired polynomials:
(a) Real targets: $a_{d}(\lambda)=2+3\lambda+\lambda^{2}$ ($\lambda_{1}=-1,\lambda_{2}=-2$) gives
$L_{o}$ $=$ $\begin{bmatrix}1.75+0.25\mathbf{j}\\ 2+0.5\mathbf{j}\end{bmatrix}$,
$L$ $=$ $\begin{bmatrix}1.25-2.25\mathbf{j}\\ -0.75\mathbf{i}+0.25\mathbf{k}\end{bmatrix}$, (26)
and the observer error dynamics matrix
$A_{\rm obs}$ $=$ $\begin{bmatrix}-2.75-\mathbf{j}&2.5\mathbf{i}-1.25\mathbf{k}\\ 0.5\mathbf{i}+0.75\mathbf{k}&-0.25-\mathbf{j}\end{bmatrix}$ (27)
has the finite right spectrum $\sigma_{\mathscr{r}}(A_{\rm obs})\!=\!\{-1,\!-2\}$ as desired.
(b) Complex pair: $a_{d}(\lambda)=2+2\lambda+\lambda^{2}$ ($\lambda_{1,2}=-1\pm\mathbf{i}$) yields
$L_{o}$ $=$ $\begin{bmatrix}1.75+0.25\mathbf{j}\\ 1+0.5\mathbf{j}\end{bmatrix}$,
$L$ $=$ $\begin{bmatrix}0.75-2.75\mathbf{j}\\ -0.25\mathbf{i}+1.75\mathbf{k}\end{bmatrix}$,
and
$A_{\rm obs}$ $=$ $\begin{bmatrix}-3.25-0.5\mathbf{j}&3\mathbf{i}-0.75\mathbf{k}\\ 2\mathbf{i}+0.25\mathbf{k}&1.25-0.5\mathbf{j}\end{bmatrix}$.
Here $\{-1\!+\!\mathbf{i},-1\!-\!\mathbf{i}\}\subset\sigma_{\mathscr{r}}(A_{\mathrm{obs}})$; equivalently, the entire spherical (similarity) class $\left[-1+\mathbf{i}\right]$ is assigned as expected.
(c) Quaternionic coefficients: Demanding specific quaternion poles (e.g., $\lambda_{1}=-1+\mathbf{j}$ and $\lambda_{2}=-2+\mathbf{k}$) sets (rounded)
$a_{d}(\lambda)$ $=$ $2.7-\mathbf{i}-1.3\mathbf{j}+0.33\mathbf{k}$
$+(3-0.67\mathbf{i}-0.33\mathbf{j}-0.33\mathbf{k})\lambda+\lambda^{2}.$
Here $a_{d}(\lambda)\in\mathbb{H}[\lambda]$ has nonreal (noncentral) coefficients, so a coordinate-free evaluation $a_{d}(A)$ is not similarity-equivariant, and we obtain (rounded)
$L_{o}$ $=$ ${\begin{bmatrix}2.42-\mathbf{i}-1.08\mathbf{j}+0.33\mathbf{k}\\ 2-0.67\mathbf{i}+0.17\mathbf{j}-0.33\mathbf{k}\end{bmatrix}}$,
$L$ $=$ ${\begin{bmatrix}-1.25-1.167\mathbf{i}-3.75\mathbf{j}-1.83\mathbf{k}\\ -1.5+1.42\mathbf{i}-1.17\mathbf{j}+1.75\mathbf{k}\end{bmatrix}}$,
and (rounded)
$A_{\rm obs}=$
$\quad\;\tfrac{1}{2}{\begin{bmatrix}\!-\!4.2\!-\!1.8\mathbf{i}\!+\!1.5\mathbf{j}\!+\!1.2\mathbf{k}&\!-\!1.8\!+\!4\mathbf{i}\!-\!1.2\mathbf{j}\!+\!1.2\mathbf{k}\\ \!-\!1.2\!+\!2\mathbf{i}\!+\!1.5\mathbf{j}\!-\!1.4\mathbf{k}&1.2\!+\!1.2\mathbf{i}\!+\!1.2\mathbf{j}\!+\!1.5\mathbf{k}\end{bmatrix}}$.
Its right spectrum is $\sigma_{\mathscr{r}}(A_{\mathrm{obs}})=\left[-1+\mathbf{i}\right]\cup[-2+\mathbf{i}]=\left[-1+\mathbf{j}\right]\cup\left[-2+\mathbf{k}\right]$ so $1\!+\!\mathbf{j},-2\!+\!\mathbf{k}\in\sigma_{\mathscr{r}}(A_{\mathrm{obs}})$ confirming that the observable-form coefficient update still assigns the desired similarity classes even when the coefficients are quaternionic.
VIII OBSERVER GAIN VIA ACKERMANN’S FORMULA
Having established the quaternionic observer in its canonical (observable) form, we now turn to a compact algebraic method for directly computing the observer gain. The resulting Observer Ackermann’s formula is the dual counterpart of the control-side expression. It provides a one–shot computation of (L) achieving any admissible placement of right–eigenvalue classes (observer poles). Its structure mirrors the classical real/complex case. Yet, its justification over $\mathbb{H}$ requires special care: the argument avoids determinants and adjugates and instead exploits the observable companion relations and the right–polynomial calculus developed earlier. This formulation establishes the quaternionic duality between controllability and observability in a determinant–free setting.
Theorem 6 (Quaternionic Observer Ackermann’s formula).
Let $A\in\mathbb{H}^{n\times n}$, $C\in\mathbb{H}^{1\times n}$ be such that the SISO pair $(A,C)$ is observable and the observability matrix $\mathcal{O}$ from (3) is invertible. Let $a_{d}(\lambda)=d_{0}+d_{1}\lambda+\cdots+d_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{R}[\lambda]$ be the desired monic error-dynamics polynomial with real coefficients, and let $e_{n}=[0,0,\dots,1]^{\mathsf{T}}\in\mathbb{R}^{n}$. Then the observer gain column
$L$ $=$ $a_{d}(A)\mathcal{O}^{-1}e_{n},$ (28)
used in (22) yields the error-dynamics matrix $A_{\mathrm{obs}}=A-LC$ whose right spectrum equals the similarity classes induced by the right roots of $a_{d}(\lambda)$:
$\sigma_{\mathscr{r}}(A-LC)$ $=$ $\mathbin{{\cup}}_{i=1}^{n}[\lambda_{i}],$ (29)
where $\{\lambda_{i}\}_{i=1}^{n}$ are the right roots of $a_{d}$ (listed with multiplicity).
Remark 7.
The restriction $a_{d}\in\mathbb{R}[\lambda]$ is essential: real coefficients lie in the center $Z(\mathbb{H})=\mathbb{R}$, so $a_{d}(A)$ is unambiguous and the required polynomial–similarity interchange is valid. The assigned classes are real or complex-conjugate spherical: for an $n$-th order SISO design, each real class consumes one degree and each nonreal class two, so $2s+r\leq n$ when placing $s$ nonreal and $r$ real classes.
Proof:
Observable companion coordinates. Choose $T$ invertible with
$A_{o}$ $=$ $T^{-1}AT,\qquad C_{o}=CT=e_{n}^{\mathsf{T}}.$
Then $A_{o}$ is in observable companion form with the rightmost column $a_{0},\ldots,-a_{n-1}]^{\mathsf{T}}$ and companion polynomial $a(\lambda)=a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}$. Let $L_{o}:=T^{-1}L$. We have
$T^{-1}(A-LC)T$ $=$ $A_{o}-L_{o}C_{o}=A_{o}-L_{o}e_{n}^{\mathsf{T}}.$ (30)
Coefficient matching in the right observable form. With $u_{k}:=A_{o}^{k}e_{1}=e_{k+1}$ for $k=0,\ldots,n-1$ and $A_{o}^{n}e_{1}=-[\,a_{0},\ldots,a_{n-1}\,]^{\mathsf{T}}$ (rightmost column read-off), we obtain
$a_{d}(A_{o})e_{1}$ $=$ $\sum_{k=0}^{n-1}d_{k}A_{o}^{k}e_{1}+A_{o}^{n}e_{1}$ (31)
$=$ $\begin{bmatrix}d_{0}\\ \vdots\\ d_{n-1}\end{bmatrix}-\begin{bmatrix}a_{0}\\ \vdots\\ a_{n-1}\end{bmatrix}=\begin{bmatrix}d_{0}-a_{0}\\ \vdots\\ d_{n-1}-a_{n-1}\end{bmatrix}$.
Setting $L_{o}:=a_{d}(A_{o})e_{1}$ therefore adds $L_{o}$ to the entries $\{-a_{k}\}$ in the rightmost column of $A_{o}$, producing the updated coefficients $\{-(a_{k}+\ell_{k})\}$ and thus the closed-loop polynomial $a_{\mathrm{obs}}(\lambda)=a_{d}(\lambda)$. Since similarity preserves the right spectrum, the error matrix ALCA-LC attains the desired right-eigenvalue classes induced by $a_{d}$.
Coordinate-free expression via $\mathcal{O}$ and central polynomials. Let $\mathcal{O}_{o}$ be the observability matrix of $(A_{o},C_{o})$. Then
$\mathcal{O}=\begin{bmatrix}C\\ CA\\ \vdots\\ CA^{n-1}\end{bmatrix}=\begin{bmatrix}C_{o}\\ C_{o}A_{o}\\ \vdots\\ C_{o}A_{o}^{n-1}\end{bmatrix}T^{-1}=\mathcal{O}_{o}T^{-1},$
so $\mathcal{O}^{-1}=T\mathcal{O}_{o}^{-1}.$ Since $u_{k}:=A_{o}^{k}e_{1}$ satisfies $u_{k}=e_{k+1}$ for $k=0,\ldots,n-1$, the first column of $\mathcal{O}_{o}$ is $e_{n}$, hence $\mathcal{O}_{o}e_{1}=e_{n}$ and $\mathcal{O}_{o}^{-1}e_{n}=e_{1}$. Therefore
$\mathcal{O}^{-1}e_{n}=Te_{1}.$
Because $a_{d}\in\mathbb{R}[\lambda]$ and $A=TA_{o}T^{-1}$,
$a_{d}(A)=T\,a_{d}(A_{o})\,T^{-1},$
and combining with $L=TL_{o}$ and $L_{o}=a_{d}(A_{o})e_{1}$ gives
$L=T\,a_{d}(A_{o})e_{1}=a_{d}(A)\,Te_{1}=a_{d}(A)\,\mathcal{O}^{-1}e_{n},$
as claimed.
Similarity preserves the right spectrum, hence $\sigma_{\mathscr{r}}(A-LC)=\sigma_{\mathscr{r}}(A_{o}-L_{o}e_{n}^{\mathsf{T}})$, and this equals the union of the similarity classes of the right zeros of $a_{d}$. ∎
Example 4 (Example 3 continued).
Building on the previous example, recall $\mathcal{O}^{-1}$ from (10). We now consider real targets only: $a_{d}(\lambda)=2+3\lambda+\lambda^{2}$ ($\lambda_{1}=-1,\lambda_{2}=-2$) gives
$a_{d}(A)=\begin{bmatrix}0.625+0.5\mathbf{j}&0.5\mathbf{i}-0.125\mathbf{k}\\ 0.5\mathbf{i}+0.125\mathbf{k}&0.625-0.5\mathbf{j}\end{bmatrix}$,
and the observer gain Ackermann formula (28) yields
$L$ $=$ $\begin{bmatrix}1.25-2.25\mathbf{j}\\ -0.75\mathbf{i}+0.25\mathbf{k}\end{bmatrix}$, (32)
which equals (26). The resulting observer error dynamics matrix $A_{\rm obs}=A-LC$ is identical to (27) and its right spectrum is $\sigma_{\mathscr{r}}(A_{\rm obs})=\{-1,-2\}$ as desired.
For complex pair targets, the results coincide with Example 3 (b). By contrast, for quaternionic targets as in Example 3 (c), the Ackermann formula returns an incorrect gain $L$ and misplaces the poles. Coefficient update in companion coordinates can handle noncentral coefficients; the Ackermann shortcut cannot.
IX SIMULATION EXAMPLE
Example 5 (based on previous examples).
Consider the SISO continuous-time quaternionic system (2) with $A$ and $C$ from (9), $B=\begin{bmatrix}1&\!\mathbf{j}\end{bmatrix}^{\!\mathsf{T}}$, and D=0D\!=\!0. Using the designs in Examples 3 and 4, we obtain the output-injection gain $L$ (see (26), (32)) that places the observer-error poles at $\lambda_{1}=-1$ and $\lambda_{2}=-2$.
With a step input $u(t)=1-\mathbf{i}+2\mathbf{j}-2\mathbf{k}$, initial states $x_{0}=\bigl[-1+\mathbf{i}-2\mathbf{j}+3\mathbf{k},\;1+2\mathbf{i}-\mathbf{j}-2\mathbf{k}\bigr]^{\mathsf{T}}$, and $\hat{x}_{0}=\bigl[0,0\bigr]^{\mathsf{T}}$, the simulation shows $\hat{x}(t)$ converging to $x(t)$ in all quaternion components, consistent with two stable real poles; see Fig. 1.
Figure 1: Observer states converge to system states in all quaternionic components.
X CONCLUSIONS
We developed a characteristic-polynomial-free observer framework for quaternionic systems, centered on an observable companion realization that handles general quaternionic targets, and a quaternionic Ackermann formula restricted to real-coefficient targets. The method enables direct placement of right-eigenvalue similarity classes in the error dynamics and clarifies controller–observer duality over $\mathbb{H},$ with the companion-coordinate coefficient update accommodating noncentral coefficients, whereas Ackermann does not. Future work will extend the approach to multi-input and discrete-time settings.

### v3

@@
 We develop observer design over hypercomplex quaternions in a characteristic-polynomial-free framework.
-Using the standard right-module convention, we derive a right observable companion form and its companion polynomial that encodes error dynamics via right-eigenvalue similarity classes.
-The design mirrors the real/complex case—coefficient updates in companion coordinates, followed by a similarity back—yet avoids determinants, characteristic/minimal polynomials, and Cayley–Hamilton identities that do not transfer to quaternions.
-We also give an Ackermann-type construction for the important case of closed-loop companion polynomials with real coefficients, ensuring similarity-equivariant evaluation.
-The results yield simple recipes for full-order observers directly over quaternions, clarify the role of right spectra and their similarity classes, and pinpoint when classical one-shot formulas remain valid.
-Numerical examples illustrate the method and advantages over vectorized or complex-adjoint surrogates.
+Using the standard right-module convention, we derive a right observable companion form and companion polynomial that encode error dynamics through right-eigenvalue similarity classes.
+We also give an Ackermann-type formula for real-coefficient target polynomials, where polynomial evaluation is similarity-equivariant.
+The resulting recipes place observer poles directly over quaternions and clarify when companion-coordinate updates and one-shot Ackermann formulas remain valid.
 I INTRODUCTION
@@
 We construct the observable companion (observability-form) realization directly from the observability matrix, link its right spectrum to companion-polynomial zeros, and derive a quaternionic Observer Ackermann formula that assigns right–eigenvalue similarity classes using a desired real-coefficient polynomial.
-Throughout, we remain in the hypercomplex (scalar) model of $\mathbb{H}$ where signals, states, and operators live directly in $\mathbb{H}.$ We work with the right spectra, and use left substitution for the companion polynomial—choices that make noncommutativity explicit and usable.
-In aerospace and robotics applications, by contrast, quaternions are typically treated in their real $4$-vector (scalar–vector) form, and observers/filters are implemented with purely real-vector operations and models [13].
-These methods are highly effective for attitude estimation and remain the default in those domains.
-Here, however, we adopt the hypercomplex definition $q\in\mathbb{H}$ and work directly with quaternionic states and operators.
-This preserves the noncommutative algebra essential for right/left spectra and similarity classes, for companion/observability forms, and for a central (real-coefficient) polynomial calculus $a(A)$ without coordinate artifacts.
-The viewpoint is also motivated by areas where the hypercomplex structure is intrinsic—e.g., quantum dynamics and control [5, 12], NMR/MRI pulse design [6], polarization/bivariate signal processing and color imaging [10], quaternion-valued filters [20], quantum computing [14] and emerging quaternionic learning models [19, 15].
+Throughout, we remain in the hypercomplex (scalar) model of $\mathbb{H}$ where signals, states, and operators live directly in $\mathbb{H}$.
+We work with the right spectrum and use left substitution for the companion polynomial—choices that make noncommutativity explicit and usable.
+From the application viewpoint, quaternion-based models appear in several areas, but are often handled after conversion to real $4$-vector coordinates, especially in aerospace and robotics, where effective real-vector observers and filters are standard [13].
+The purpose of the present paper is different: we show that observer design can also be developed directly over $\mathbb{H}$ in a way that preserves the native hypercomplex structure.
+This matters precisely in settings where one wants to reason in terms of quaternionic spectra, similarity classes, companion forms, and central polynomial evaluation rather than through coordinate embeddings.
+Such a viewpoint is natural not only in quaternionic system theory itself, but also in areas where the hypercomplex structure is intrinsic or algorithmically meaningful, including quantum dynamics and control [5, 12], NMR/MRI pulse design [6], polarization/bivariate signal processing and color imaging [10], quaternion-valued filtering [22], quantum computing [14], and emerging quaternionic learning models [21, 15].
+Application-oriented quaternionic observer constructions for four-channel systems are discussed in [20].
 We believe the observer results established here over $\mathbb{H}$—including the observable companion construction and the quaternionic Observer Ackermann formula—are original.
@@
 With $u\!\equiv\!0$, stability means $x(t)\!\to\!0$ (CT) or $x(k)\!\to\!0$ (DT) for all initial states.
-Equivalently, using the standard right eigenvalues $\lambda$ of $A$, the system is stable iff $\operatorname{Re}\lambda<0$ for all $\lambda$; (see [16]).
+Equivalently, using the standard right eigenvalues $\lambda$ of $A$, the system is stable iff $\operatorname{Re}\lambda<0$ for all $\lambda$; see [16].
 Controllability
@@
   2. 2.
-    Hence, the right spectrum of $A_{o}$ is exactly the union of the similarity classes of the right zeros of $a(\lambda)$:
-    $\sigma_{\mathscr{r}}(A_{o})=\mathbin{{\cup}}_{i}[\lambda_{i}],$
-    counting multiplicities.
+    Hence, the right spectrum of $A_{o}$ consists exactly of the similarity classes of the right zeros of $a(\lambda)$.
   3. 3.
@@
   2. 2.
-    Hence, the right spectrum of $A$ is exactly the union of the similarity classes of the right zeros of $a(\lambda)$: $\sigma_{\mathscr{r}}(A)=\mathbin{{\cup}}_{i}[\lambda_{i}]$, counting multiplicities.
+    Hence, the right spectrum of $A$ consists exactly of the similarity classes of the right zeros of $a(\lambda)$.
   3. 3.
@@
 Different cyclic choices of $C$ may yield different polynomials.
-Still, their right zeros agree up to quaternionic similarity classes: for every such $C$, the right zeros of the associated polynomial form the same multiset of classes—the right spectrum of $A$.
+Still, their right zeros agree up to quaternionic similarity classes: for every such $C$, the right zeros of the associated polynomial determine the same set of similarity classes, namely the right spectrum of $A$.
 For a fixed pair $(A,C)$, however, the observable companion $(A_{o},C_{o})$ and the attached polynomial are unique.
@@
 It enables implementable state feedback $u=K\hat{x}$, monitoring, and prediction.
-The relations below are standard from the real/complex literature and will be referenced later.
-They also hold for quaternionic systems over $\mathbb{H}$ with minor modifications that are detailed in a later section.
+For quaternionic systems, the same observer structure is used, but pole assignment must be interpreted in terms of right-eigenvalue similarity classes.
 A well-known Luenberger full-order observer [11] is described by
@@
 VI DUALITY OVER $\mathbb{H}$
-Over $\mathbb{R}$, transpose drives controller–observer duality: controllability of $(A,B)$ is equivalent to observability of $(A^{\mathsf{T}},B^{\mathsf{T}})$, and one often maps gains via $L=K^{\mathsf{T}}$ (cf.
-[2]).
-Over $\mathbb{H}$, this fails because the correct adjoint is the conjugate transpose $M^{*}=\overline{M}^{\mathsf{T}}$, which alone preserves the quaternionic Hermitian form and converts left to right eigenpairs with eigenvalue conjugation [21, 18, 7, 1].
-Thus, duality must be formulated with and with explicit care for right/left eigenpairs (see also quaternion-valued system treatments [9, 4]).
-The original system and its dual system are related through mappings
-$A\leftrightarrow A^{*},B\leftrightarrow C^{*},C\leftrightarrow B^{*},D\leftrightarrow D^{*}.$
-Their controllability and observability are related via
+Over $\mathbb{R}$, controller–observer duality is expressed through transpose: controllability of $(A,B)$ is equivalent to observability of $(A^{\mathsf{T}},B^{\mathsf{T}})$, and one typically transfers gains via $L=K^{\mathsf{T}}$; see, e.g., [2].
+Over $\mathbb{H}$, the transpose is no longer the correct dual object.
+The appropriate operation is the conjugate transpose $M^{*}=\overline{M}^{\mathsf{T}}$, since it preserves the quaternionic Hermitian form, $\langle Ax,y\rangle=\langle x,A^{*}y\rangle,$ and is compatible with quaternionic spectral theory [23, 18, 7, 1].
+Thus the duality mapping must be written as
+$A\leftrightarrow A^{*},\qquad B\leftrightarrow C^{*},\qquad C\leftrightarrow B^{*},\qquad D\leftrightarrow D^{*}.$
+Accordingly, the original system and its dual are linked by
 $\mathcal{O}_{n}(A^{*},B^{*})$ $=$ $\mathcal{C}_{n}(A,B)^{*},$
-see [21, 18], and therefore
+because $(AB)^{*}=B^{*}A^{*}$; see [23, 18]. Since rank is preserved under , this yields
 $(A,B)\ \text{controllable}\ \Longleftrightarrow\ (A^{*},B^{*})\ \text{observable},$
 $(A,C)\ \text{observable}\ \Longleftrightarrow\ (A^{*},C^{*})\ \text{controllable}.$
-If $K\in\mathbb{H}^{1\times n}$ is a controller for the dual pair $(A^{*},C^{*})$, then the observer (output-injection) gain for the original system is
+Therefore, if $K\in\mathbb{H}^{1\times n}$ is a state-feedback gain for the controllable dual pair $(A^{*},C^{*})$, then the corresponding observer gain for the original system is
 $L$ $=$ $K^{*}\in\mathbb{H}^{n\times 1}.$
-The error and dual closed-loop matrices satisfy $\sigma_{\mathscr{r}}\big(A-LC\big)=\sigma_{\mathscr{r}}\big(A^{*}-C^{*}K\big).$ Since $q]=[\overline{q}]$ for quaternionic similarity classes, the two spectra coincide as class sets.
-See [7, 1].
-Proof sketches and references
-The transpose fails over $\mathbb{H}$; only the adjoint satisfies $\langle Ax,y\rangle=\langle x,A^{*}y\rangle$ and aligns with quaternionic spectral theory [21, 18, 7, 1].
-The dual mapping $(A,B,C,D)\!\leftrightarrow\!(A^{*},C^{*},B^{*},D^{*})$ is standard.
-Using $(AB)^{*}=B^{*}A^{*}$ and rank invariance yields $\mathcal{O}_{n}(A^{*},B^{*})=\mathcal{C}_{n}(A,B)^{*}$, hence $(A,B)$ controllable $\Leftrightarrow$ $(A^{*},B^{*})$ observable and $(A,C)$ observable $\Leftrightarrow$ $(A^{*},C^{*})$ controllable [21, 18].
-Design $K$ on $(A^{*},C^{*})$ and set $L=K^{*}$; then $\sigma_{\mathscr{r}}(A-LC)=\overline{\sigma_{\mathscr{r}}(A^{*}-C^{*}K)}=\sigma_{\mathscr{r}}(A^{*}-C^{*}K)$ by conjugation invariance of right-eigenvalue classes [7, 1].
+Moreover,
+$\sigma_{\mathscr{r}}(A-LC)=\sigma_{\mathscr{r}}(A^{*}-C^{*}K),$
+because the adjoint converts right eigenvalues into conjugates, and quaternionic similarity classes satisfy $q]=[\overline{q}]$; see [7, 1].
+Hence both matrices have the same assigned right-eigenvalue similarity classes.
+See also quaternion-valued system treatments in [9, 4].
 VII QUATERNIONIC OBSERVER POLE PLACEMENT VIA COMPANION FORM
@@
 there exists a unique observer gain $L\in\mathbb{H}^{n\times 1}$ such that the right spectrum of
-$A_{\mathrm{obs}}$ $:=$ $A-LC$ (24)
-equals the multiset of right zeros (similarity classes, with multiplicities) of $a_{d}$. Moreover, with $L_{o}:=T^{-1}L$ we have
+$A_{\mathrm{obs}}:=A-LC$
+consists exactly of the similarity classes induced by the right roots of $a_{d}$, counted with multiplicity. Moreover, with $L_{o}:=T^{-1}L$ we have
 $L$ $=$ $TL_{o},$ (24)
@@
 $A_{\rm obs}$ $=$ $\begin{bmatrix}-2.75-\mathbf{j}&2.5\mathbf{i}-1.25\mathbf{k}\\ 0.5\mathbf{i}+0.75\mathbf{k}&-0.25-\mathbf{j}\end{bmatrix}$ (26)
-has the finite right spectrum $\sigma_{\mathscr{r}}(A_{\rm obs})\!=\!\{-1,\!-2\}$ as desired.
+has right spectrum $\sigma_{\mathscr{r}}(A_{\mathrm{obs}})=\{[-1],[-2]\}$, with both classes singleton.
 (b) Complex pair: $a_{d}(\lambda)=2+2\lambda+\lambda^{2}$ ($\lambda_{1,2}=-1\pm\mathbf{i}$) yields
@@
 $A_{\rm obs}$ $=$ $\begin{bmatrix}-3.25-0.5\mathbf{j}&3\mathbf{i}-0.75\mathbf{k}\\ 2\mathbf{i}+0.25\mathbf{k}&1.25-0.5\mathbf{j}\end{bmatrix}$.
-Here $\{-1\!+\!\mathbf{i},-1\!-\!\mathbf{i}\}\subset\sigma_{\mathscr{r}}(A_{\mathrm{obs}})$; equivalently, the entire spherical (similarity) class $\left[-1+\mathbf{i}\right]$ is assigned as expected.
-(c) Quaternionic coefficients: Demanding specific quaternion poles (e.g., $\lambda_{1}=-1+\mathbf{j}$ and $\lambda_{2}=-2+\mathbf{k}$) sets (rounded)
+Here the assigned class is $\left[-1+\mathbf{i}\right]\in\sigma_{\mathscr{r}}(A_{\mathrm{obs}})$; equivalently, its standard representatives are $1\pm\mathbf{i}$.
+(c) Quaternionic coefficients: Choosing target similarity classes represented, for example, by $\lambda_{1}=-1+\mathbf{j}$ and $\lambda_{2}=-2+\mathbf{k},$ sets (rounded)
 $a_{d}(\lambda)$ $=$ $2.7-\mathbf{i}-1.3\mathbf{j}+0.33\mathbf{k}$
@@
 $\quad\;\tfrac{1}{2}{\begin{bmatrix}\!-\!4.2\!-\!1.8\mathbf{i}\!+\!1.5\mathbf{j}\!+\!1.2\mathbf{k}&\!-\!1.8\!+\!4\mathbf{i}\!-\!1.2\mathbf{j}\!+\!1.2\mathbf{k}\\ \!-\!1.2\!+\!2\mathbf{i}\!+\!1.5\mathbf{j}\!-\!1.4\mathbf{k}&1.2\!+\!1.2\mathbf{i}\!+\!1.2\mathbf{j}\!+\!1.5\mathbf{k}\end{bmatrix}}$.
-Its right spectrum is $\sigma_{\mathscr{r}}(A_{\mathrm{obs}})=\left[-1+\mathbf{i}\right]\cup[-2+\mathbf{i}]=\left[-1+\mathbf{j}\right]\cup\left[-2+\mathbf{k}\right]$ so $1\!+\!\mathbf{j},-2\!+\!\mathbf{k}\in\sigma_{\mathscr{r}}(A_{\mathrm{obs}})$ confirming that the observable-form coefficient update still assigns the desired similarity classes even when the coefficients are quaternionic.
+Its right spectrum consists of the two similarity classes
+$\sigma_{\mathscr{r}}(A_{\mathrm{obs}})=\{[-1+\mathbf{i}],[-2+\mathbf{i}]\}=\{[-1+\mathbf{j}],[-2+\mathbf{k}]\}.$
+Thus the observable-form coefficient update still assigns the desired similarity classes even when the coefficients are quaternionic.
 VIII OBSERVER GAIN VIA ACKERMANN’S FORMULA
@@
 The resulting Observer Ackermann’s formula is the dual counterpart of the control-side expression.
-It provides a one–shot computation of (L) achieving any admissible placement of right–eigenvalue classes (observer poles).
+It provides a one–shot computation of $L$ achieving any admissible placement of right–eigenvalue classes (observer poles).
 Its structure mirrors the classical real/complex case.
@@
 Theorem 6 (Quaternionic Observer Ackermann’s formula).
-Let $A\in\mathbb{H}^{n\times n}$, $C\in\mathbb{H}^{1\times n}$ be such that the SISO pair $(A,C)$ is observable and the observability matrix $\mathcal{O}$ from (3) is invertible.
-Let $a_{d}(\lambda)=d_{0}+d_{1}\lambda+\cdots+d_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{R}[\lambda]$ be the desired monic error-dynamics polynomial with real coefficients, and let $e_{n}=[0,0,\dots,1]^{\mathsf{T}}\in\mathbb{R}^{n}$.
-Then the observer gain column
+Let $A\in\mathbb{H}^{n\times n}$, $C\in\mathbb{H}^{1\times n}$ be such that the SISO pair $(A,C)$ is observable and the observability matrix $\mathcal{O}$ from (3) is invertible. Let
+$a_{d}(\lambda)=d_{0}+d_{1}\lambda+\cdots+d_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{R}[\lambda]$
+be the desired monic error-dynamics polynomial with real coefficients, and let $e_{n}=[0,0,\dots,1]^{\mathsf{T}}\in\mathbb{R}^{n}$. Then the observer gain column
 $L$ $=$ $a_{d}(A)\mathcal{O}^{-1}e_{n},$ (27)
 used in (22) yields the error-dynamics matrix $A_{\mathrm{obs}}=A-LC$ whose right spectrum equals the similarity classes induced by the right roots of $a_{d}(\lambda)$:
-$\sigma_{\mathscr{r}}(A-LC)$ $=$ $\mathbin{{\cup}}_{i=1}^{n}[\lambda_{i}],$ (29)
-where $\{\lambda_{i}\}_{i=1}^{n}$ are the right roots of $a_{d}$ (listed with multiplicity).
+$\sigma_{\mathscr{r}}(A-LC)$ $=$ $\{[\lambda_{i}]:i=1,\ldots,n\},$ (28)
+where $\{\lambda_{i}\}_{i=1}^{n}$ are right roots of $a_{d}$ (listed with multiplicity).
 Remark 7.
-The restriction $a_{d}\in\mathbb{R}[\lambda]$ is essential: real coefficients lie in the center $Z(\mathbb{H})=\mathbb{R}$, so $a_{d}(A)$ is unambiguous and the required polynomial–similarity interchange is valid.
+The restriction
+$a_{d}\in\mathbb{R}[\lambda]$
+is essential: real coefficients lie in the center $Z(\mathbb{H})=\mathbb{R}$, so $a_{d}(A)$ is unambiguous and the required polynomial–similarity interchange is valid.
 The assigned classes are real or complex-conjugate spherical: for an $n$-th order SISO design, each real class consumes one degree and each nonreal class two, so $2s+r\leq n$ when placing $s$ nonreal and $r$ real classes.
@@
 as claimed.
-Similarity preserves the right spectrum, hence $\sigma_{\mathscr{r}}(A-LC)=\sigma_{\mathscr{r}}(A_{o}-L_{o}e_{n}^{\mathsf{T}})$, and this equals the union of the similarity classes of the right zeros of $a_{d}$.
+Similarity preserves the right spectrum, hence $\sigma_{\mathscr{r}}(A-LC)=\sigma_{\mathscr{r}}(A_{o}-L_{o}e_{n}^{\mathsf{T}})$, and this spectrum consists exactly of the similarity classes of the right zeros of $a_{d}$.
 ∎
@@
 $L$ $=$ $\begin{bmatrix}1.25-2.25\mathbf{j}\\ -0.75\mathbf{i}+0.25\mathbf{k}\end{bmatrix}$, (31)
-which equals (26). The resulting observer error dynamics matrix $A_{\rm obs}=A-LC$ is identical to (27) and its right spectrum is $\sigma_{\mathscr{r}}(A_{\rm obs})=\{-1,-2\}$ as desired.
+which equals (25).
+The resulting observer error dynamics matrix $A_{\rm obs}=A-LC$ is identical to (26) and its right spectrum is $\sigma_{\mathscr{r}}(A_{\mathrm{obs}})=\{[-1],[-2]\}$, as desired, with both classes singleton.
 For complex pair targets, the results coincide with Example 3 (b).
@@
 Example 5 (based on previous examples).
-Consider the SISO continuous-time quaternionic system (2) with $A$ and $C$ from (9), $B=\begin{bmatrix}1&\!\mathbf{j}\end{bmatrix}^{\!\mathsf{T}}$, and D=0D\!=\!0.
-Using the designs in Examples 3 and 4, we obtain the output-injection gain $L$ (see (26), (32)) that places the observer-error poles at $\lambda_{1}=-1$ and $\lambda_{2}=-2$.
-With a step input $u(t)=1-\mathbf{i}+2\mathbf{j}-2\mathbf{k}$, initial states $x_{0}=\bigl[-1+\mathbf{i}-2\mathbf{j}+3\mathbf{k},\;1+2\mathbf{i}-\mathbf{j}-2\mathbf{k}\bigr]^{\mathsf{T}}$, and $\hat{x}_{0}=\bigl[0,0\bigr]^{\mathsf{T}}$, the simulation shows $\hat{x}(t)$ converging to $x(t)$ in all quaternion components, consistent with two stable real poles; see Fig.
+For (2) with $A,C$ from (9), $B=\begin{bmatrix}1&\!\mathbf{j}\end{bmatrix}^{\!\mathsf{T}}$, $D=0$, input $u(t)=1-\mathbf{i}+2\mathbf{j}-2\mathbf{k}$, and initial conditions $x_{0}=\bigl[-1+\mathbf{i}-2\mathbf{j}+3\mathbf{k},\;1+2\mathbf{i}-\mathbf{j}-2\mathbf{k}\bigr]^{\mathsf{T}}$, $\hat{x}_{0}=\bigl[0,0\bigr]^{\mathsf{T}}$, the gain from (25)–(31) yields convergence $\hat{x}(t)\to x(t)$ in all quaternionic components; see Fig.
 1.
@@
 X CONCLUSIONS
-We developed a characteristic-polynomial-free observer framework for quaternionic systems, centered on an observable companion realization that handles general quaternionic targets, and a quaternionic Ackermann formula restricted to real-coefficient targets.
-The method enables direct placement of right-eigenvalue similarity classes in the error dynamics and clarifies controller–observer duality over $\mathbb{H},$ with the companion-coordinate coefficient update accommodating noncentral coefficients, whereas Ackermann does not.
-Future work will extend the approach to multi-input and discrete-time settings.
+We developed a characteristic-polynomial-free observer framework for quaternionic systems, centered on an observable companion realization and a quaternionic Ackermann formula for real-coefficient targets.
+The method assigns right-eigenvalue similarity classes directly in the error dynamics and makes the -based controller–observer duality explicit; companion-coordinate updates accommodate noncentral coefficients, whereas Ackermann does not.
+Application-oriented quaternionic observers for four-channel systems are discussed in [20], while related controllability companion-form and state-feedback results are developed in [19].